OpenStudy (anonymous):

Solve for x: 3x − 24 = 81 fifty-seven over three one hundred five over three 57 105

OpenStudy (anonymous):

The answer is 35: 3x-24=81 3x=81+24 NOW: divide both sides by 3. x=105/3 x=35